OpenStudy (anonymous):

a family travelled a total distance of 740 km in 12 hrs and 15 mins on a trip. part of the trip was by car at 80 km/hr. and the remainder by ferry at 16 km/hr. how many hrs. were spent travelling by car?

OpenStudy (anonymous):

where are you so far?

OpenStudy (cwrw238):

here are some equations arising from the problem c + f = 740 80 = c/Tc 16 = f/Tf Tc + Tf = 12.25 where c = distance travelled by car f = ,,,,,,,,,,,,,,,,,,,,,,,,,,,,by ferry Tc = time by car Tf = ...........ferry

OpenStudy (cwrw238):

aim to get 2 equations in Tc and Tf

OpenStudy (anonymous):

Seeing the time of taking car is x hrs. \[80x+16(12\frac{ 1 }{ 4 }-x)=740\] \[80x+16\times12\frac{ 1 }{ 4 }-16x=740\] \[64x+196=740\] \[64x=740-196\] \[64x=544\] \[x=544\div64\] \[x=8.5\] The time that travelling by car was 8.5 hours.